What mass of HF must be used to produce 1.5L of an aqueous solution with a pH of 2.00? 


1
Expert's answer
2021-05-24T07:20:04-0400

Solution:

We can convert between pH and [H3O+] using the following equation:

pH = - log[H3O+]

[H3O+] = 10-pH

[H3O+] = 10-2.00 = 0.01

[H3O+] = 0.01M


HF is weak acid.

The acid dissociation reaction for HF is the following:

HF(aq) + H2O(l) ⇌ F-(aq) + H3O+(aq)

Ka for HF = 6.6×10-4 at 25oC

The Ka expression for HF is:




ICE Table: 



αC = [H3O+] = [F-] = 0.01 M

Hence,

[HF] = [H3O+] × [F-] / Ka

[HF] = (0.01 × 0.01) / (6.6×10-4) = 0.15 M

C(HF) = [HF] + αC = 0.15 + 0.01 = 0.16 M

C(HF) = 0.16 M


n(HF) = C(HF) × V

n(HF) = 0.16 M × 1.5 L = 0.24 mol

The molar mass of HF is 20.01 g/mol

Hence,

(0.24 mol HF) × (20.01 g HF / 1 mol HF) = 4.8024 g HF = 4.8 g HF


Answer: 4.8 grams of HF must be used.